OpenStudy (anonymous):

PLEASE HELP!! Describe the properties of logs used to solve the following equation. Please also provide the solution. log2 (4) + log2 (x+6) = 2log2 (6)

OpenStudy (anonymous):

do you know your log rules/properties?

OpenStudy (anonymous):

not really, i'm confused with it.

OpenStudy (anonymous):

ok, there actually three main ones:

OpenStudy (anonymous):

use the product rule

OpenStudy (anonymous):

there's the product rule, quotient rule, and power rule

OpenStudy (anonymous):

i know that theres product property equality property

OpenStudy (anonymous):

log_b (XY) = log_b X + log_b Y log_b (X/Y) = log_b X - log_b Y log_b (X^Y) = Y*log_b X

OpenStudy (anonymous):

these are the 3...

OpenStudy (lgbasallote):

there was this guy that did a tutorial on these stuffs here on openstudy..wait i'll look for it...maybe it can help you

OpenStudy (anonymous):

thanks @lgbasallote , :)

OpenStudy (anonymous):

this is your equation: log2 (4) + log2 (x+6) = 2log2 (6) notice we can use the first rule on that left side to get: log2 [(4)(x+6)] = 2log2 (6)

OpenStudy (anonymous):

this one right log_b (XY) = log_b X + log_b Y

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now the right side...

OpenStudy (lgbasallote):

aww darn -_- i cant access my old questions darn lag :/ sorry cant give the tutorial link haha

OpenStudy (anonymous):

soright... we can manage... @lgbasallote , thanks though.... :) now the right side: log2 [(4)(x+6)] = log2 (6^2)

OpenStudy (anonymous):

do you know what property i used on the right side?

OpenStudy (anonymous):

the same one right.

OpenStudy (anonymous):

no, I used the third rule i wrote down...

OpenStudy (anonymous):

This has all the propertys of logs if you need them: http://www.andrews.edu/~calkins/math/webtexts/numb17.htm#BAS

OpenStudy (lgbasallote):

well just for future references... http://openstudy.com/study#/updates/4fa45f8fe4b029e9dc34e0b5

OpenStudy (anonymous):

okay and that's the final answer?

OpenStudy (anonymous):

no, you need to solve the equation...

OpenStudy (anonymous):

okay so it looks like this right now. log2 [(4)(x+6)] = 2log2 (6) and log2 [(4)(x+6)]= log2 (6^2)

OpenStudy (anonymous):

you'll need to solve this equation in order to get your answer: (4)(x+6)]= (6^2) 4x + 24 = 36

OpenStudy (anonymous):

okay and i know that we got rid of the log2 because all of them are the same

OpenStudy (anonymous):

correct... can you tell me the answer then?

OpenStudy (anonymous):

need help?